Circular No. 8350

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 K3 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to have cometary appearance on CCD images taken by P. Birtwhistle (Great Shefford, U.K.; coma diameter 12′′ with a hint of a 9′′-wide extension emanating from a 3′′ central condensation for ≈ 20′′ in p.a. 90^{o}) and by P. Holvercem and M. Schwartz (Tenagra Observatories, Nogales, AZ; 0.81-m f/7 reflector; 5′′ coma).

2004 UT  α_{2000}  δ_{2000}  Mag.
May 29.34561  20^{h}39^{m}47.^{s}43  +38^{o}05′13′′.0  19.3

The early astrometry, preliminary parabolic orbital elements (T = 2004 June 30.0 TT, q = 1.106 AU, i = 112^{o}.0, ω = 60^{o}.2, Ω = 275^{o}.6, equinox 2000.0), and an ephemeris appear on MPEC 2004-L04.

SUPERNOVAE 2004bz, 2004ca, 2004cb, 2004cc Several apparent supernovae have been discovered on unfiltered CCD

images — 2004cb by L. A. G. Monard, Pretoria, South Africa (0.30-m reflector) and the rest by the LOSS team (via W. Li; KAIT detector).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004bz June 2.46 22^{h}15^{m}01.^{s}05 +15^{o}24′16′′.6 18.3 32′′.2 W, 12′′.3 N
2004ca June 2.47 21 43 32.84 +43 41 33.0 18.2 24′′.8 W, 6′′.2 S
2004cb May 12.80 13 43 48.52 −29 44 59.2 15.8 10′′ W, 21′′ S
2004cc June 10.26 12 36 34.40 +11 14 32.8 17.5 1′′.9 E, 12′′.8 N

Additional magnitudes by the respective discoverers: SN 2004bz in MCG +02-56-25, 2003 Nov. 29.17 UT, [19.5; 2004 June 3.48, 18.2. SN 2004ca in UGC 11799, May 25.45, [19.0; June 3.48, 18.0. SN 2004cb in an edge-on galaxy, Apr. 16.814, [18.5; June 7.758, 15.8 ± 0.3. SN 2004cc in NGC 4568, May 23.26, [19.0; June 11.23, 17.1. SN 2004cb is also absent on the red Digitized Sky Survey field (limiting mag 20).

Further to IAUC 8338, T. Matheson et al. report that a spectrum (range 340–730 nm) of SN 2004ca, obtained by M. Calkins on June 10.46 UT, shows it to be a type-Ia supernova near maximum spectral-feature age 2 ± 2 days before maximum light). Adopting the NED recession velocity of 5338 km/s for the host galaxy, the supernova expansion velocity is 12800 km/s for Si II (rest 635.5 nm) and 20400 km/s for Ca II (rest 395.1 nm).

2004 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green